# Large Language Models Share Representations of Latent Grammatical Concepts Across Typologically Diverse Languages

### v1

@@ 4.1 Experimental Setup
This is an adaptation and translation of data from Arora et al. (2024). Then, we check whether the probability of the grammatical sequence is higher for each pair. In our experiments, we only consider sentences where Llama-3-8B is indeed capable of making grammatically correct predictions.
@@
In our setting, we follow Templeton et al. (2024) and intervene in the model by scaling a single SAE feature to a multiple of its maximum activation value observed in the dataset. [^7] For this purpose, we record the maximum activations of features across 20 M tokens. More formally, for a given grammatical concept, we select one feature from the set of multilingual features identified in §3 and scale it to a multiple of the observed activation value. We implement feature steering by decomposing the residual stream activations ${\mathbf{x}}$ into the sum of two components, the SAE reconstruction and the reconstruction error (see Equation 1). Then, we replace the SAE term with a modified SAE reconstruction in which we set the activation value of a specific feature in the SAE to a counterfactual value, but leave the error term unchanged. We then run the forward pass on the network using this modified residual stream. We only apply the intervention to the last token position of each generation step to prevent degenerate outputs. Thus, for a given morphosyntactic concept, similar to Templeton et al. (2024), we explore different multiples and choose the one that works best empirically.
We consider activating features in contexts where they would not be active otherwise. We observe that intervening on multiple features at the same time is very sensitive to the combination of features and their activation value. Therefore, we follow Templeton et al. (2024) in intervening on only a single feature at a time. To select the feature, we manually inspect the features and select one feature per concept where the activation values are most interpretable.
@@ 4.2 Quantitative Results
We define the efficacy of the intervention as the proportion of examples where the probe’s label flips after the intervention. To measure this, we generate a translation with and without the intervention. Then, in new forward passes with no interventions, we give the model these translations, pool their representations, and measure whether the probe’s (from §2) labels differ between the two translations. We find that intervening on a single feature is often sufficient to change the models prediction towards the indended concept (e.g., a masculine instead of feminine pronoun; see Figure 5). However, we also observe that, for some concepts, intervening on a single feature is often insufficient.
Figure 6: Fraction of times an intervention on a grammatical concept exclusively impacts that concept, indicating the degree of cross-concept interference. We observe that interference is generally very rare.
@@
We also evaluate the selectivity of the interventions. Specifically, we want interventions to only flip the labels on the concept that we intervene on, and not others. We operationalise this using the probing classifiers trained in §2. As before, we generate two translations: one with and without feature interventions to a given concept. Then, we compute selectivity as the proportion of translation pairs for which none of the other concept probes changed their predictions. Our results (Figure 6) suggest that the interventions are generally selective to the concept.
@@ 4.3 Qualitative Results
The quantitative results provide insights into the generalisation of the identified features to a different distribution, and allow us to quantify the success rate of the interventions. However, the dataset is neither naturalistic nor diverse. Therefore, we perform the same interventions on the Flores-101 dataset, which provides aligned translations across 101 languages (Goyal et al., 2021). We present a selection of qualitative examples in Figure 7.
However, we acknowledge that the intervention was not successful in some cases, and sometimes led to degenerate generations. Concept-specific tuning over the number of features and activation values used was also required. These limitations highlight common problems with feature steering methods: activations often must be scaled to significantly larger values than would naturally be observed in practice (Templeton et al., 2024). This, combined with intervening on multiple features simultaeously, can result in unexpected or degenerate outputs (Shabalin et al., 2023).
@@ 5 Related Work
A central question in multilingual language modeling is whether language models develop a universal concept representation, disentangled from specific languages. Huang et al. (2023) studied different forms of cross-lingual transfer using a variety of models and probing tasks. Wendler et al. (2024) found that languages models trained on unbalanced, English-dominated corpora operate in a concept space that lies closer to English than to other languages. Ferrando (2024) studies the specific mechanism of subject-verb agreement and finds that these mechanisms are consistent across languages. Dumas et al. (2024) observed that concepts (e.g. “car”) are encoded disentangled from specific languages and can be transferred between them. Similarly, Stanczak et al. (2022) found that probes trained on different languages read the same concepts from the same neurons in smaller-scale masked language models such as mBERT (Devlin et al., 2019) and XLM-R (Conneau et al., 2020), while Feng et al. (2024) observed that their propositional probes generalise to Spanish. De Varda and Marelli (2023) extended this study by investigating the cross-lingual consistency of individual neurons responding to syntactic phenomena. Our work extends these findings to large-scale models not explicitly intended to be multilingual, and shows stronger causal evidence for the functional selectivity of these conceptual representations.
@@ 6 Discussion and Conclusion
Language models trained primarily on English data perform surprisingly well in other languages. Prior work has emphasized the importance of balanced pretraining corpora, so as to not overfit to a single language (Conneau et al., 2020; Mohammadshahi et al., 2022); however, more recent studies suggest that the size of the pretraining corpus—not balance—may be more important for cross-lingual generalization (Jiang et al., 2023; Dubey et al., 2024; Schäfer et al., 2024). Why is this effective? Our results provide evidence that large-scale pretraining, even with imbalanced corpora, induces equally cross-lingualistcally generalizable grammatical abstracts as pretraining with balanced corpora. This is evidence in favor of a hypothesis discussed in Wendler et al. (2024): the internal lingua franca of large language models may not be English words per se, but rather concepts. That said, it is still reasonable to assume these concepts are likely biased toward how English handles them in models pre-trained on imbalanced corpora.
@@
Sparse autoencoders are a method for unsupervised search for features. However, they will only recover a feature if it is linearly encoded. Thus, if a concept is encoded in a non-linear manner, then sparse features may struggle to recover the full range of that variable. Thus, sparse features are generally best-suited to encoding binary relations, whereas some features may be encoded in more complex arrangements such as circular shapes (cf. Engels et al. 2024).
Feature interventions.
We use SAE features to intervene on the models internal state. Similar to Templeton et al. (2024), we find that this typically requires clamping the feature activations to values outside their observed range over the training dataset. However, clamping feature activations to extreme values often causes the model to produce nonsensical generations, e.g., repeating the same token indefinitely. Thus, while our results demonstrate that the identified features can be causally responsible for model behavior, optimally intervening on groups of features, is an open problem that future work should address.
Distinction between model understanding and generation.
In our experiments, the linear probes measure whether it is possible to linearly separate grammatical concepts in the model’s representations, whereas the machine translation experiments measure the effect of feature directions on model generation. Past work (Meng et al., 2022; Orgad et al., 2024; Gottesman and Geva, 2024) has observed a distinction between a model encoding a concept versus being able to use that concept during generation (i.e., a “knowing vs. saying” distinction). Anecdotally, we observed that some features with a high causal effect on the linear probe do not necessarily have a similar effect on the model’s generation. We hypothesize that this is due to an intrinsic distinction between concept understanding/representation and concept generation.
Human biases in feature interpretation.
One must be cautious in applying human concepts to language model representations. It is probable that language models deploy distinct concept spaces from humans; thus, human explanations of neurons or features are likely to be biased in ways that can result in suboptimal predictions of when the neuron or feature will activate (cf. Huang et al., 2023).
@@ A.2 Libraries
To extract activations and intervene on the model, we use nnsight (Fiotto-Kaufman et al., 2024). For training the linear probes, we use scikit-learn (Pedregosa et al., 2011). For training the SAEs, we use dictionary-learning (Marks and Mueller, 2024).
@@ Appendix C Attribution Patching
$\hat{\text{IE}}_{\text{ig}}(m;{\mathbf{a}};x_{\text{clean}},x_{\text{patch}})=$
$\left(\sum_{\alpha}\nabla_{{\mathbf{a}}}m\big{|}_{\alpha{\mathbf{a}}_{\text{clean}}+(1-\alpha){\mathbf{a}}_{\text{patch}}}\right)({\mathbf{a}}_{\text{patch}}-{\mathbf{a}}_{\text{clean}})$
where the sum ranges over $K=10$ equally-equally spaced $\alpha\in\{0,\tfrac{1}{K},\dots,\tfrac{K-1}{K}\}$. While this cannot be done in parallel for two nodes when one is downstream of another, it can be done in parallel for arbitrarily many nodes which do not depend on each other. Thus the additional cost of computing $\hat{\text{IE}}_{\text{ig}}$ over $\hat{\text{IE}}_{\text{atp}}$ scales linearly in $K$.
@@ D.1 Gated Sparse Autoencoders
The SAE architecture, as introduced in Bricken et al. (2023) is defined by encoder weights $W_{e}\in\mathbb{R}^{m\times n}$, decoder weights $W_{d}\in\mathbb{R}^{n\times m}$ with columns constrained to have a $L_{2}$-norm of 1, and biases $b_{e}\in\mathbb{R}^{m}$, $b_{d}\in\mathbb{R}^{n}$. Given an input $\mathbf{x}\in\mathbb{R}^{n}$, the SAE computes
@@ D.3 Performance Across Languages
We evaluate the performance of the sparse autoencoder trained on Llama-3-8B using the multilingual C4 (mC4) dataset (Xue et al., 2021), itself a multilingual extension of the C4 dataset (Raffel et al., 2020). Table 2 presents the loss recovered when replacing the residual stream activations with the sparse autoencoder reconstructions across a range of languages. For this purpose, we compute the loss recovered as
$L_{recovered}=\frac{L_{reconstructed}-L_{zero}}{L_{original}-L_{zero}}$ (7)
where $L_{original}$ is the cross entropy (CE) loss of the language model without intervention, $L_{zero}$ is the CE loss of the model when zero ablating the activations, and $L_{reconstructed}$ is the loss when replacing the activations with the sparse autoencoder reconstructions. This is a standard extrinsic metric to measure autoencoder quality (Mueller et al., 2024). However, it does not measure the intrinsic interpretability of the SAE features, which is not possible to measure in the absence of ground truth features (Karvonen et al., 2024; Makelov et al., 2024).
@@ Appendix E Features Across Languages
Appendix F Machine Translation
@@ F.1 Prompt Format
The following is an example of our 2-shot translation prompt:
@@
F.2 Translation Performance
We evaluate the ability of Llama-3-8B and Aya-23-8B to translate languages in-context using the Flores-101 dataset (Goyal et al., 2021). Specifically, we measure the BLEU score (Papineni et al., 2002) using the SacreBLEU implementation (Post, 2018). We prompt models with 2 ground-truth exemplar translations, and then give it the test sentence in the source language (see Figure 14). Results are presented in Figure 15 and 16.
@@ F.2 Translation Performance
Appendix G Counterfactual Dataset

### v2

@@ 4.1 Experimental Setup
This is an adaptation and translation of data from Arora et al. (2024). Then, we check whether the probability of the grammatical sequence is higher for each pair. In our experiments, we only consider sentences where Llama-3-8B is capable of making grammatically correct predictions. We find that it is almost always correct.
@@
We follow Templeton et al. (2024) and intervene in the model by setting a single SAE feature to a multiple of its maximum activation value observed in the dataset. [^7] For this purpose, we record the maximum activations of each feature across 20M tokens from mC4 (Xue et al., 2021). Specifically, we decompose the residual stream activations ${\mathbf{x}}$ into the sum of two components, the SAE reconstruction and the reconstruction error (see Eq. 1). Then, we replace the SAE term with a modified SAE reconstruction in which we set the activation value of a specific feature to a counterfactual value, but leave the error term unchanged. We then run the forward pass on the network using this modified residual stream. We only apply the intervention to the last token position of each generation step to prevent degenerate outputs.
To select the feature to intervene with, we manually inspect feature activation patterns. We select one feature per concept from the set of multilingual features identified in §3 where the activation values are most interpretable. Similar to Templeton et al. (2024), we explore different multiples and choose the one that works best empirically.
@@ 4.2 Quantitative Results
We define the efficacy of the intervention as the proportion of examples where the probe’s label flips after the intervention. To measure this, we generate a translation with and without the intervention. Then, in new forward passes with no interventions, we give the model these translations, pool their representations, and measure whether the probe’s (from §3.1) labels differ between the two translations. We find that intervening on a single feature is often sufficient to change the models prediction towards the indended concept (e.g., a masculine instead of feminine pronoun; see Figure 5). However, for some concepts, intervening on a single feature is often insufficient.
Figure 6: Fraction of times an intervention on a grammatical concept exclusively impacts that concept, indicating the degree of cross-concept interference. We find that interference is generally very rare. Note that Turkish does not inflect for grammatical gender.
@@
We also evaluate the selectivity of the interventions. Specifically, we want interventions to only flip the labels on the concept that we intervene on, and not others. We operationalise this using the probing classifiers trained in §3.1. As before, we generate two translations: one with and without feature interventions to a given concept. Then, we compute selectivity as the proportion of translation pairs for which none of the other concept probes changed their predictions. Our results (Figure 6) suggest that the interventions are generally selective to the concept.
@@ 4.3 Qualitative Results
The quantitative results provide insights into the generalisation of the identified features to a different distribution, and allow us to quantify the success rate of the interventions. However, the dataset is neither naturalistic nor diverse. Therefore, we perform the same interventions on the Flores-101 dataset (Goyal et al., 2022), which provides aligned translations across 101 languages. We present a selection of qualitative examples in Figure 7.
However, we acknowledge that the intervention was not successful in some cases, and sometimes led to degenerate generations. Concept- and language-specific tuning over the activation values was also required. These limitations highlight common problems with feature steering methods, including that activations must often be scaled to significantly larger values than would naturally be observed (Templeton et al., 2024). This, combined with intervening on multiple features simultaneously, can result in unexpected or degenerate outputs (Shabalin et al., 2023).
@@ 5 Related Work
A central question in multilingual language modeling is whether LMs develop universal concept representations, disentangled from specific languages. Dhar and Bisazza (2021) study different forms of cross-lingual transfer in recurrent neural networks and found that exposing our LMs to a related language does not always increase gram- matical knowledge in the target language. Ruzzetti et al. (2023) study the similarity of monolingual encoder models, finding higher similarity between models trained for typologically similar languages. Wendler et al. (2024) find that languages models trained on unbalanced, English-dominated corpora operate in a concept space that lies closer to English than to other languages. Ferrando (2024) find that mechanisms of subject-verb agreement are consistent across languages. Dumas et al. (2024) observe that concepts (e.g., “car”) are encoded disentangled from specific languages and can be transferred between them. Similarly, Stanczak et al. (2022) find that probes trained on different languages read the same concepts from the same neurons in smaller-scale masked language models such as mBERT (Devlin et al., 2019) and XLM-R (Conneau et al., 2020), while Feng et al. (2024) observe that their propositional probes generalise to Spanish. de Varda and Marelli (2023) extend this study by investigating the cross-lingual consistency of individual neurons responding to syntactic phenomena. Our work extends these findings to large-scale models (including one not explicitly intended to be multilingual), and shows stronger causal evidence for the functional selectivity of these conceptual representations.
@@ 6 Discussion and Conclusion
Language models trained primarily on English data perform surprisingly well in other languages. Prior work has emphasized the importance of balanced pretraining corpora, so as to not overfit to a single language (Conneau et al., 2020; Mohammadshahi et al., 2022); however, more recent studies suggest that the size of the pretraining corpus—not balance—may be more important for cross-lingual generalization (Jiang et al., 2023; Dubey et al., 2024; Schäfer et al., ). Why is this effective? Our results provide evidence that large-scale pretraining, even with imbalanced corpora, induces equally cross-lingualistically generalizable grammatical abstractions as pretraining with balanced corpora. This is evidence in favor of a hypothesis discussed in Wendler et al. (2024): the internal lingua franca of large language models may not be English words per se, but rather concepts. That said, it is still reasonable to assume these concepts are likely biased toward how English handles them in models pre-trained on imbalanced corpora.
@@
Steering with SAEs.
We use SAE features to intervene on the models internal state. Similar to Templeton et al. (2024), we find that steering with SAEs typically requires clamping feature activations to values outside their observed range over the training dataset. However, clamping feature activations to extreme values often causes the model to produce nonsensical generations, e.g., repeating the same token indefinitely. SAEs have also been shown to perform worse for fine-grained steering model behavior than alternatives (Wu et al., 2025), such as difference-in-means (Marks and Tegmark, 2024). Thus, while our results demonstrate that the identified features can be causally responsible for model behavior, optimally steering model behavior is an open problem that future work should address.
Distinction between model understanding and generation.
In our experiments, the linear probes measure whether it is possible to linearly separate grammatical concepts in the model’s representations, whereas the machine translation experiments measure the effect of feature directions on model generation. Past work (Meng et al., 2022; Orgad et al., 2025; Gottesman and Geva, 2024) has observed a distinction between a model encoding a concept versus being able to use that concept during generation (i.e., a “knowing vs. saying” distinction). Anecdotally, we observed that some features with a high causal effect on the linear probe do not necessarily have a similar effect on the model’s generation. We hypothesize that this is due to an intrinsic distinction between concept understanding/representation and concept generation.
Human biases in feature interpretation.
One must be cautious in applying human concepts to LM representations. It is probable that LMs deploy distinct concept spaces from humans; thus, human explanations of neurons or features are likely to be biased in ways that can result in suboptimal predictions of when the neuron or feature will activate (cf. Huang et al., 2023).
@@
Sparse autoencoders are a method for unsupervised search for features. However, they will only recover a feature if it is linearly encoded. Thus, if a concept is encoded in a non-linear manner, then sparse features may struggle to recover the full range of that variable. Thus, sparse features are generally best-suited to encoding binary relations, whereas some features may be encoded in more complex arrangements such as circular shapes (cf. Engels et al. 2025; Csordás et al. 2024).
Acknowledgements
Jannik Brinkmann is supported by the German Federal Ministry for Economic Affairs and a grant from the Long-Term Future Fund. Chris Wendler is supported by a grant from Open Philanthropy. Aaron Mueller is supported by a postdoctoral fellowship under the Zuckerman STEM Leadership Program.
@@ A.2 Libraries
To extract activations and intervene on the model, we use nnsight (Fiotto-Kaufman et al., 2025). For training the linear probes, we use scikit-learn (Pedregosa et al., 2011). For training the SAEs, we use dictionary-learning (Marks and Mueller, 2024), itself based on nnsight.
@@ Appendix C Attribution Patching
$\hat{\text{IE}}_{\text{ig}}(m;{\mathbf{a}}_{\text{clean}},{\mathbf{a}}_{\text{patch}})=$ (3)
$\left(\sum_{\alpha}\left.\nabla_{{\mathbf{a}}}m\right|_{\alpha{\mathbf{a}}_{\text{clean}}+(1-\alpha){\mathbf{a}}_{\text{patch}}}\right)({\mathbf{a}}_{\text{patch}}-{\mathbf{a}}_{\text{clean}})$
where the sum in (2) ranges over $K=10$ equally-equally spaced $\alpha\in\{0,\tfrac{1}{K},\dots,\tfrac{K-1}{K}\}$. While this cannot be done in parallel for two nodes when one is downstream of another, it can be done in parallel for arbitrarily many nodes which do not depend on each other. Thus the additional cost of computing $\hat{\text{IE}}_{\text{ig}}$ over $\hat{\text{IE}}_{\text{atp}}$ scales linearly in $K$.
@@ D.1 Gated Sparse Autoencoders
The Gated SAE architecture, as introduced in Bricken et al. (2023) is defined by encoder weights $W_{e}\in\mathbb{R}^{m\times n}$, decoder weights $W_{d}\in\mathbb{R}^{n\times m}$ with columns constrained to have a $L_{2}$-norm of 1, and biases $b_{e}\in\mathbb{R}^{m}$, $b_{d}\in\mathbb{R}^{n}$. Given an input $\mathbf{x}\in\mathbb{R}^{n}$, the SAE computes
@@ D.3 Performance Across Languages
We evaluate the performance of the SAE trained on Llama-3-8B using the multilingual C4 (mC4) dataset (Xue et al., 2021), itself a multilingual extension of the C4 dataset (Raffel et al., 2020). Table 2 presents the loss recovered when replacing the residual stream activations with the SAE reconstructions across a range of languages. For this purpose, we compute the loss recovered as
$\mathcal{L}_{\text{recovered}}=\frac{\mathcal{L}_{\text{reconstructed}}-\mathcal{L}_{\text{zero}}}{\mathcal{L}_{\text{original}}-\mathcal{L}_{\text{zero}}}$ (8)
where $\mathcal{L}_{\text{original}}$ is the cross entropy (CE) loss of the language model without intervention, $\mathcal{L}_{\text{zero}}$ is the CE loss of the model when zero ablating the activations, and $\mathcal{L}_{\text{reconstructed}}$ is the loss when replacing the activations with the SAE reconstructions. This is a standard extrinsic metric to measure autoencoder quality (Mueller et al., 2024). However, it does not measure the intrinsic interpretability of the SAE features, which is not possible to measure in the absence of ground truth features (Karvonen et al., 2024; Makelov et al., 2024).
@@ Appendix E Features Across Languages
Appendix F Machine Translation Details
@@ F.1 Prompt Format
We present an example of our 2-shot translation prompt in Figure 14.
F.2 Translation Performance
We evaluate the ability of Llama-3-8B and Aya-23-8B to translate languages in-context using the Flores-101 dataset (Goyal et al., 2022). Specifically, we measure the BLEU score (Papineni et al., 2002) using the SacreBLEU implementation (Post, 2018). We prompt models with 2 ground-truth exemplar translations, and then give it the test sentence in the source language (see Figure 14). Results are presented in Figure 15 and 16.
@@ F.2 Translation Performance
F.3 Directional Translation Results
In Fig. 5, we present averages for language pairs translating in both directions (e.g., English to French and French to English). Here, we provide additional details about the results for each translation direction separately (see Fig. 17).
@@ F.3 Directional Translation Results
Turkish Hemşire sonraki adımlardan emin değil
Hemşireler sonraki adımlardan emin değiller
@@
Turkish Hasan arıyor çünkü o hastadır
Hasan aradı çünkü o hastaydı